OpenStudy (anonymous):

Evaluate lim x͢ 0 x/1-√1-x

ganeshie8 (ganeshie8):

\[\large \lim \limits_{x \to 0} ~\dfrac{x}{1-\sqrt{1-x}}\]

ganeshie8 (ganeshie8):

like this ?

ganeshie8 (ganeshie8):

heard of the term `conjugate` before ?